OpenStudy (anonymous):

Medal?!?! Can someone tell me what a linear function is and give an example on how to do it? Please be specific when explaining.

OpenStudy (anonymous):

Definition: A function f is linear if it can be expressed in the form f ( x) = mx + b where m and b are constants and x is an arbitrary member of the domain of f. Often the relationship between two variables x and y is a linear function expressed as an equation y = mx + b . Examples: Let v be the downward velocity of a body falling freely near the surface of the Moon under the influence of gravity and let t be the time it has fallen. Let v0 be the velocity imparted to it initially. Then 0 v = gt + v . Thus, v is a linear function of t with m = g, the acceleration due to gravity near the surface of the Moon, and b = v0. ) Suppose that you put P dollars into a savings account that draws simple interest at a rate of r. After time t, the value of the savings account will be V(t) = P (1+ rt). V is a linear function of t. In this example, it is not quite so obvious what the constants m and b are. We leave it as an exercise to identify them. With solving this might help. https://www.khanacademy.org/math/cc-eighth-grade-math/cc-8th-relationships-functions/analyzing-functions-algebra/v/interpreting-features-of-linear-functions-example

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Welcome, and let me know if you need any more :) GOOD LUCK BEAUTIFUL!